Circular No. 5963

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994I IN NGC 5194 M. P. Rupen and R. A. Sramek, National Radio Astronomy Observa-

tory; S. D. Van Dyk, University of California at Berkeley; K. W. Weiler, Naval Research Laboratory; and N. Panagia, Space Telescope Science Institute, report: “Observations on Apr. 3.17 UT with the Very Large Array in its highest-resolution (A) configuration at wavelengths 1.3 and 3.6 cm (22.5, 8.4 GHz) show an unresolved (<0′′.3) radio source within 2′′ of the optical position of the supernova (IAUC 5961). Flux densities were 2.32 ± 0.52 mJy at 1.3 cm and 0.55 ± 0.07 mJy at 3.6 cm, where the uncertainties include both random and systematic (calibration) errors; the random noise in these images is 0.23 and 0.04 mJy/beam, respectively. By contrast, SN 1993J was still undetected at radio wavelengths even three days after optical discovery (the first radio detection—0.75 mJy at 1.3 cm—occurred five days afterward; cf. IAUC 5743, 5752). Since SN 1994I is more distant than SN 1993J, this difference in time of turn-on is intrinsic and probably means that the envelope of circumstellar material around SN 1994I is relatively thin. However, the strong radio emission does suggest a sufficiently dense circumstellar medium that many of the associated effects sometimes seen at other wavelengths (x-ray/ultraviolet emission, narrow optical lines, infrared dust echoes, etc.) should also be apparent in SN 1994I. We find a position for the supernova of α = 13^{h}27^{m}47^{s}.731, δ = +47^{o}26′57′′.84 (equinox 1950.0; uncertainty ± 0′′.1 in each coordinate); we also detect radio emission from the nucleus of M51, giving an offset for the SN of 14′′.3 east and 12′′.3 south. VLA observations from 1982 to 1983 (Ford et al. 1985, Ap.J. 293, 132) at 6 and 20 cm give for the nucleus of M51 the same position (to within 0′′.13) and consistent flux densities, but show no source at the position of the supernova (6-cm flux density < 0.1 mJy). Radio monitoring is continuing at these and other wavelengths.”

M. W. Richmond, Princeton University; and A. V. Filippenko, R. R. Treffers, and S. D. Van Dyk, University of California at Berkeley, report that NGC 5194 has been monitored as part of the Leuschner Observatory Supernova Search; SN 1994I was not present to a limiting R magnitude of ∼ 16.2 on Mar. 28 UT. Preliminary reductions of data from the LOSS 0.76-m reflector and from the 0.5-m Berkeley Automatic Imaging Telescope yield the following: Apr. 2.5, B = 14.38, V = 14.00, R = 13.87, I = 13.82; 3.4, V = 13.54, R = 13.43, I = 13.55; 4.4, U = 13.79, B = 13.97, V = 13.38, R = 13.20, I = 13.19.

1994 April 4 Daniel W. E. Green